OpenStudy (hkmiu):

Can you help me find the inverse of y = -3x+5 step by step?

OpenStudy (anonymous):

Sure :) Can u inverse 6 ?

OpenStudy (anonymous):

@hkmiu :Please answer to my questions I want to u learn step by step :)

OpenStudy (anonymous):

huh?

OpenStudy (hkmiu):

Oh sorry, I was doing my other math problems!

OpenStudy (hkmiu):

what do you mean by inverse 6?

OpenStudy (anonymous):

No problem :) inverse 6 ... It s an example to start :)

OpenStudy (hkmiu):

1/6?

OpenStudy (anonymous):

yep ! Thank u Modom :) Now1/ x ...

OpenStudy (anonymous):

huh?

OpenStudy (hkmiu):

inverse of 1/x ???????

OpenStudy (hkmiu):

x/1 ??????????? idk ? >.<

OpenStudy (anonymous):

yep ! Why u think dont know u was ritht :) and now -3x ?

OpenStudy (anonymous):

huh ? dont afreid and answer Modom :)

OpenStudy (hkmiu):

1/-3x?

OpenStudy (anonymous):

Thank u Modom :) and now : 5 ?

OpenStudy (hkmiu):

1/5

OpenStudy (anonymous):

thank u again :) and now : y ?

OpenStudy (hkmiu):

1/y ????

OpenStudy (anonymous):

yep ! Now come back to your question ! You answered to your self :)

OpenStudy (hkmiu):

Wait how?!

OpenStudy (anonymous):

We had :y = -3x+5 u find the inverse of y ... Then 5 than -3x Just it :D

OpenStudy (anonymous):

Got it ?:)

OpenStudy (anonymous):

inverse of y ? u said 1/y inverse of 5? u said 1/5 inverse of -3x ?u said 1/-3x and just it and end the answer we have inverse of y = -3x+5___>1/y=1/-3x+1/5 Understoond ?:)

OpenStudy (hkmiu):

That's not one of the answer my teacher gave me!

OpenStudy (hkmiu):

:O

OpenStudy (anonymous):

u can answer this from all the users :D

OpenStudy (anonymous):

u st ill need help

OpenStudy (hkmiu):

Yess ;(

OpenStudy (anonymous):

hey fan me and ill do the same.. ill help tho

OpenStudy (hkmiu):

Is it (5 - x)/3?

OpenStudy (skullpatrol):

Nope, define what an "inverse" is first please?

OpenStudy (skullpatrol):

The question asks you to find the inverse, so you should first have a clear idea of what you are looking for, right?

OpenStudy (hkmiu):

I might.

OpenStudy (hkmiu):

I tried working the problem with the same step as my teacher's video but it doesnt seems right

OpenStudy (skullpatrol):

Start with the definition.

OpenStudy (hkmiu):

opposite in direction

OpenStudy (skullpatrol):

So, how does that relate to this question?

OpenStudy (hkmiu):

find the opposite of this equation

OpenStudy (skullpatrol):

What would that mean?

OpenStudy (hkmiu):

......i don't know........

OpenStudy (skullpatrol):

With this equation if someone gave us values of x we could output values of y, right?

OpenStudy (hkmiu):

Yeah, switch the x and y

OpenStudy (hkmiu):

\[x = -3y + 5\]

OpenStudy (skullpatrol):

Why did you do that?

OpenStudy (hkmiu):

beauseee thats the first step to finding the inverse of a linear function ... at least thats what the teacher told me and all the yt videos told me...

OpenStudy (skullpatrol):

OK, now what do you do?

OpenStudy (hkmiu):

ummm, subtract both sides by 5

OpenStudy (hkmiu):

\[x-5=-3y\]

OpenStudy (skullpatrol):

OK, now...???

OpenStudy (hkmiu):

divide bot sides by -3 \[\left( x-5 \right)/-3 = y\]

OpenStudy (skullpatrol):

Is that the answer your teacher gave?

OpenStudy (hkmiu):

nope....

OpenStudy (anonymous):

ok i see that skullpatrol.. is doing a fine job.. so ll scoot ;-)!

OpenStudy (skullpatrol):

Try just working with y = -3x + 5 and solve it for x =?

OpenStudy (hkmiu):

-3x+8 would be (8-x)/3 right??

OpenStudy (hkmiu):

wait no the signs are wrong.

OpenStudy (skullpatrol):

Nope work with y = -3x + 5 and solve it for x =?

OpenStudy (hkmiu):

(x-8)/-3?

OpenStudy (hkmiu):

uhhhh

OpenStudy (skullpatrol):

work with y = -3x + 5 and solve it for x =?

OpenStudy (hkmiu):

(y-5)/-3

OpenStudy (skullpatrol):

What happened to "x"?

OpenStudy (hkmiu):

it just flip flopped.

OpenStudy (hkmiu):

well x = (y-5)/-3

OpenStudy (skullpatrol):

$$\huge x=\huge -\frac{y-5}{3}$$

OpenStudy (skullpatrol):

That^ is the inverse.

OpenStudy (hkmiu):

Sooo the equation y= -3x+8 would be (x-8)/-3?

OpenStudy (skullpatrol):

Is that the answer your teacher gives?

OpenStudy (hkmiu):

This one is a free response so i wouldnt know ><

OpenStudy (hkmiu):

waittt why would it be x = (y-5)/3 ?? why isnt the 3 negative

OpenStudy (anonymous):